OpenStudy (anonymous):

[(5-3)*5]+2*7 divided by 7

OpenStudy (jack1):

\[[(5-3) \times 5]+2 \times 7\div 7\] \[[(2) \times 5]+2 \times 7\div 7\] \[[10]+2 \times 7\div 7\] \[[10]+2 \times \frac 77\] \[[10]+2 \times \frac 11\] \[[10]+2 \times 1\] \[[10]+2 \] =...? notes: refer to the BODMAS / PEDMAS system (whichever you were taught)

OpenStudy (anonymous):

12

OpenStudy (jack1):

purrfect @countrygal2013